OpenStudy (anonymous):

I don't get this: Emily earns $48,500 per year. Her personal allowance is $6475. Her income is then taxed at 20% up to $37,400 and then at 40% above $37,400. How much does Emily earn after tax?

OpenStudy (anonymous):

Okay *-^

OpenStudy (anonymous):

Sorry I made a mistake

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

Hehe

OpenStudy (texaschic101):

Isn't your personal allowance income that is NOT taxed ? So wouldn't you take 6475 from 48,500 first ? Then figure out the taxes after that ?

OpenStudy (anonymous):

Yah that was what I thought too T-T

OpenStudy (anonymous):

So i think you subtract 6475 off 48,500 coz it is what she spends before tax is collected right?!

OpenStudy (anonymous):

But then what!??

OpenStudy (anonymous):

I would take away 6475. from 48500 and then apply the 20% to the first 37,000 ad then the 40% to the rest

OpenStudy (anonymous):

Okay then ^-^

OpenStudy (texaschic101):

48,500 - 6475 = 42,025 -- taxable income 20% up to 37,400 = 0.20(37,400) = 7480 42,025 - 37,400 = 4624 40% of 4624 = .40(4624) = 1849.60 total taxes : 7480 + 1849.60 = 9329.60 (taxes she has to pay) she would then earn : I am not sure if you add in the 6475 un-taxable income... so I am not sure if the answer is 48,500 - 9329.60 or if it is 42,025 - 9329

OpenStudy (anonymous):

I get it now!!! Thank you very much!!!! ^-^